# Camera: Multi-Matrix Joint Compression for MoE Models via Micro-Expert Redundancy Analysis

### v1

@@ 1 Introduction
We frame micro-expert pruning as the problem of removing a fixed proportion (e.g., 20%) of micro-experts while minimizing degradation in decoding performance. As we detail later, this is an NP-hard problem that cannot be solved exactly within practical time constraints. To address this challenge, we propose Camera, an efficient and accurate approximation algorithm for estimating the importance ranking of micro-experts in MoE layers. Building on this analysis, we propose Camera-P, a structured pruning framework that jointly prunes redundant micro-experts across the three FFN weight matrices, as shown in Figure 1. We further extend this idea to mixed-precision quantization by proposing Camera-Q, a novel micro-expert-aware partitioning scheme for assigning precision. Experiments show that Camera-P can prune MoE models ($>$50B) within 5 minutes on a single NVIDIA A100-40GB GPU—orders of magnitude faster than existing methods that often require hours of computation on multi-GPUs. Both Camera-P and Camera-Q consistently outperform strong baselines across nine zero-shot tasks and diverse MoE models, demonstrating superior efficiency and generalization.
@@ 2.2 Pruning Methods for MoE Models
[…]
@@ 3.2 Redundancy Problem
This is similar to a class of problems known as Column Subset Selection Problems (CSSP). Such combinatorial optimization problems have been proven to be NP-hard, making it impossible to obtain exact solutions in polynomial time (Shitov 2021). Although general approximate solutions exist (Mahoney and Drineas 2009; Ordozgoiti, Canaval, and Mozo 2018), the large $N_{e}$ still poses significant challenges for efficient hardware utilization. For instance, models like Mixtral-8$\times$7B and Deepseek-MoE-16B have $N_{e}$ on the order of $10^{5}$. Even with approximation, the time and space costs for these models remain substantial.
@@ 3.3 Camera Algorithm
[…]
@@
[…]
@@
[…]
@@ 4.1 Pruning Framework Camera-P
We aim to prune a specified proportion $\lambda$ of micro-experts in each MoE module. Leveraging the Camera algorithm, we identify the less important micro-experts in each layer. Specifically, for each selected micro-expert, all three associated weight vectors are simultaneously zeroed out. The detailed procedure is outlined in Algorithm 2, which applies pruning layer by layer. First, lines 575$\sim$ 75 ∼ 7 collect calibration samples from the MoE module, including the input and output hidden states. Then, the Camera greedy algorithm is used to rank all micro-experts and determine the set $S$ of those to retain. Within each expert, only the micro-experts with indices in $S$ are preserved. Finally, the post-pruned output $\mathbf{O}$ is recomputed, proceeding to the next year.
@@ 4.2 Experimental Setup
[…]
@@ 5.2 Experimental Setup
We perform 2-bit mixed-precision quantization on MoE layers on Deepseek-MoE-16B and Qwen3-30B-A3B, focusing on comparisons with single-precision GPTQ (Frantar et al. 2022) and the recent MC method (Huang et al. 2025) that applies mixed precision at the expert level. For fairness, Camera-Q also adopts GPTQ for its quantization step (Lines 10$\sim$12 in Algorithm 3). We further introduce a variant baseline, Camera-Q, which applies mixed precision by slicing all matrices along the input dimension (i.e., column-wise), following the most common practice. Specifically, our method is $\mathbf{W}_{s_{k},:}^{\mathrm{up}},\mathbf{W}_{s_{k},:}^{\mathrm{gate}},\mathbf{W}_{:,s_{k}}^{\mathrm{down}}$, while Camera-Q uses $\mathbf{W}_{:,s_{k}}^{\mathrm{up}},\mathbf{W}_{:,s_{k}}^{\mathrm{gate}},\mathbf{W}_{:,s_{k}}^{\mathrm{down}}$, breaking precision consistency within each micro-expert. We analyze this difference in detail in Appendix A.6. Precision settings are in Appendix A.5. All baselines are calibrated on the C4 dataset, and our method uses 128 sequences of length 2048. Evaluation follows the same setting as Camera-P.
@@ A.2 Proof Details
[…]
@@
[…]
@@
[…]
@@ A.3 Supplementary Micro-Expert Distributions
Mixtral-8$\times$7B (8 common experts, top-2 experts are activated)
@@
Table 5: Supplementary pruning results of evaluation experiment on Phi3.5-MoE-42B and Mixtral-8$\times$7B. The best scores are in bold. We also test the performance of the original model (16-bit) as a reference.
    Model Mixtral-8$\times$7B Phi3.5-MoE-42B Deepseek-MoE-16B Qwen3-30B-A3B Deepseek-V2
@@ A.4 Supplementary Pruning Experiments
In our primary experiments, we evaluate the effectiveness of the Camera-P algorithm on models featuring a larger number of smaller experts. This design choice is strongly supported by recent developments in the open-source community. Over the past two years, leading organizations and research groups—including Alibaba (Qwen series) and Deepseek—have consistently advanced the state-of-the-art by adopting MoE architectures that mix more lightweight experts. Notable examples include Qwen1.5-MoE-A2.7B, Qwen2-57B-A14B, Qwen3-30B-A3B, and the Deepseek-MoE family. In contrast, earlier MoE models often employ fewer but heavier experts, leading to higher activation ratios and lower computational efficiency, as shown in Table 6. To further assess the performance of Camera-P under such settings ($N_{E}\leq 16$), we conduct additional experiments on Phi3.5-MoE-42B and Mixtral-8$\times$7B, as shown in Table 5.
@@ A.5 Mixed-precision Settings in Camera-Q
[…]
@@ A.6 Integrity of Micro-Experts
Table 7: Comparison of intermediate hidden states after pruning from different algorithms. The experimental setup follows previous sections. For Mixtral-8$\times$7B, we prune 25% of the experts and report results for layers 1, 16, and 31; for Deepseek-MoE-16B, we prune 20% and report layers 2, 12, and 22. “L2” denotes the L2 norm of the difference between the pruned hidden states and those of the original model, while “Cosine” refers to the cosine similarity between them.
@@ A.8 Details on Baselines
  • $D^{2}$-MoE: We use the official open-source implementation. The calibration dataset is of the same source as that used in Camera-P, with the data size following the open-source recommendation (approximately 5$\sim$10x larger than that of Camera-P). The shared base weight are selected based on the Fisher criterion, with pp_ratio set to 0.1 to preserve model capacity as much as possible. The delta_ratio is determined using the search code from the author, and share_ratio is set to 1.

### v2

@@ 1 Introduction
We frame micro-expert pruning as the problem of removing a fixed proportion (e.g., 20%) of micro-experts while minimizing degradation in decoding performance. As we detail later, this is an NP-hard problem that cannot be solved exactly within practical time constraints. To address this challenge, we propose Camera, an efficient and accurate approximation algorithm for estimating the importance ranking of micro-experts in MoE layers. Building on this analysis, we propose Camera-P, a structured pruning framework that jointly prunes redundant micro-experts across the three FFN weight matrices, as shown in Figure 1. We further extend this idea to mixed-precision quantization by proposing Camera-Q, a novel micro-expert-aware partitioning scheme for assigning precision. Experiments show that Camera-P can prune MoE models (>>50B) within 5 minutes on a single NVIDIA A100-40GB GPU—orders of magnitude faster than existing methods that often require hours of computation on multi-GPUs. Both Camera-P and Camera-Q consistently outperform strong baselines across nine zero-shot tasks and diverse MoE models, demonstrating superior efficiency and generalization.
@@ 2.2 Pruning Methods for MoE Models
Most pruning methods for MoE models focus on expert-level compression rather than on individual matrices. These fall mainly into two categories: direct pruning and expert merging (sometimes followed by pruning). Direct pruning methods like NAEE exhaustively evaluate expert combinations and retain the one yielding the lowest loss on calibration data (Lu et al. 2024). While effective on Mixtral-MoE, this brute-force approach does not scale well to modern sparse MoEs. MoE-$I^{2}$ (Yang et al. 2024) and MoE-Pruner (Xie et al. 2024) partially prune expert weights, but struggle to balance identifying important weights and achieving speedup. Expert merging methods fall into two subgroups. The first assumes functional similarity across experts, grouping and merging them based on activation frequency, output similarity, or knowledge distribution (e.g., MC-SMoE (Li et al. 2024), HC-SMoE (Chen et al. 2024), TAP (Zhang et al. 2024), EEP (Liu et al. 2024b)). This idealized assumption often limits performance. The second extracts common components from all experts (e.g., $D^{2}$-MoE (Gu et al. 2025), ResMoE (Ai et al. 2025), Sub-MoE (Li et al. 2025a), MoE-SVD (Li et al. 2025b)), then applies low-rank compression to approximate residuals. However, these approaches overlook the functional integrity across the three transformations in FFNs and often fail to preserve the most critical parameters. Our method, Camera-P, belongs to the partial pruning family but differs in two key ways. First, it performs fine-grained pruning at the micro-expert level, capturing cross-matrix coordination in FFNs. Second, it is powered by Camera, our fast and accurate importance estimator, addressing a major bottleneck of existing methods.
@@ 3.2 Redundancy Problem
[…]
@@ 3.3 Camera Algorithm
Based on the definition of energy, we can rank all micro-experts and prioritize retaining the Top-$|S|$ highest-energy ones. Before presenting the specific algorithm, we first provide its tighter error bound.
@@
Let $\hat{\mathbf{Y}}$ denote the decoding result using the Top-$|S|$ highest-energy micro-experts, and let $\mathbf{Y}^{*}$ denote the rank-$|S|$ SVD approximation of $\mathbf{Y}$. If $k=N_{e}-|S|$, the approximation error of $\hat{\mathbf{Y}}$ differs from the optimal SVD by only an $O(k)$-delta, i.e.,
@@
10:[$\mathcal{E}\leftarrow[(1-\alpha)\boldsymbol{\phi}^{2}+\alpha\boldsymbol{\phi}^{\infty}]\cdot\boldsymbol{\omega}^{2}$
@@ 4.1 Pruning Framework Camera-P
We aim to prune a specified proportion $\lambda$ of micro-experts in each MoE module. Leveraging the Camera algorithm, we identify the less important micro-experts in each layer. Specifically, for each selected micro-expert, all three associated weight vectors are simultaneously zeroed out. The detailed procedure is outlined in Algorithm 2, which applies pruning layer by layer. First, lines $5\sim 7$ collect calibration samples from the MoE module, including the input and output hidden states. Then, the Camera greedy algorithm is used to rank all micro-experts and determine the set $S$ of those to retain. Within each expert, only the micro-experts with indices in $S$ are preserved. Finally, the post-pruned output $\mathbf{O}$ is recomputed, proceeding to the next year.
@@ 4.2 Experimental Setup
We compare Camera-P with two strong baselines: NAEE (Lu et al. 2024), which prunes entire experts directly, and $D^{2}$-MoE (Gu et al. 2025), a recent strong merge-then-compress approach. To address the scalability issue with $N_{E}\geq 8$ in NAEE, we follow MoE-$I^{2}$ (Yang et al. 2024) and adopt efficient genetic search. Camera-P uses $\alpha=0.95,0.95,1.00$ for the three models.
@@ 5.2 Experimental Setup
[…]
@@ A.2 Proof Details
Let $\hat{\mathbf{Y}}$ denote the decoding result using the Top-$|S|$ highest-energy micro-experts, and let $\mathbf{Y}^{*}$ denote the rank-$|S|$ SVD approximation of $\mathbf{Y}$. If $k=N_{e}-|S|$, the approximation error of $\hat{\mathbf{Y}}$ differs from the optimal SVD by only an $O(k)$-delta, i.e.,
@@
By the Eckart-Young-Mirsky theorem, we know that the truncated SVD gives the matrix with the closest distance to $\mathbf{Y}$ among all matrices whose rank does not exceed $|S|$. The approximation error is:
@@
The theorem is proved. This shows that the decoding error of the MoE layer after micro-expert pruning differs from the optimal rank‑$|S|$ approximation by at most a value $\delta\left(O(k)\right)$ proportional to $k$. The magnitude of $\delta$ depends on the energy distribution of the micro-experts: if their energies vary significantly and the boundary energy $\varepsilon$ is faint, then $\delta$ will also be small. In such cases, the energy-based pruning algorithm Camera-P can achieve a good and controllable approximation. Figure 5 confirms our assumption that the energy distribution of micro-experts is highly non-uniform, which ensures the effectiveness of Camera-P.
@@ A.3 Supplementary Micro-Expert Distributions
[…]
@@
[…]
@@ A.4 Supplementary Pruning Experiments
[…]
@@ A.5 Mixed-precision Settings in Camera-Q
In the experiments in Table 2 of the main text, MC assigns different precisions across experts, whereas Camera-Q and Camera-Q apply varying precisions across micro-experts within each expert. Consequently, the weight matrix of each expert consists of multiple precision levels. We implement Camera-Q and its variants using a three-level mixed-precision scheme. To achieve an average bit-width of $b$, we define three precision levels: [b+1,b,b1][b+1,b,b-1] bits, with their respective proportions in the total weights set to [r,12r,r][r,1-2r,r]. For example, if the desired average precision is 2 bits with $r=0.1$, then the bit-widths $[3,2,1]$ are assigned in proportions $[0.1,0.8,0.1]$, yielding an average of $3\times 0.1+2\times 0.8+1\times 0.1=2.0$ bits. Since we perform group-wise quantization over every 128 vectors, each group stores an additional zero-point and scale factor, totaling $2\times 16=32$ bits (following the GPTQ algorithm (Frantar et al. 2022)). This adds an overhead of $32/128=0.25$ bit per weight, resulting in a total average cost of 2.25 bits per weight.
@@ A.6 Integrity of Micro-Experts
[…]
@@ A.8 Details on Baselines
[…]